### v1

# COBASE: A new copula-based shuffling method for ensemble weather forecast postprocessing

Abstract
Weather predictions are often provided as ensembles generated by repeated runs of numerical weather prediction models. These forecasts typically exhibit bias and inaccurate dependence structures due to numerical and dispersion errors, requiring statistical postprocessing for improved precision. A common correction strategy is the two-step approach: first adjusting the univariate forecasts, then reconstructing the multivariate dependence. The second step is usually handled with nonparametric methods, which can underperform when historical data are limited. Parametric alternatives, such as the Gaussian Copula Approach (GCA), offer theoretical advantages but often produce poorly calibrated multivariate forecasts due to random sampling of the corrected univariate margins. In this work, we introduce COBASE, a novel copula-based postprocessing framework that preserves the flexibility of parametric modeling while mimicking the nonparametric techniques through a rank-shuffling mechanism. This design ensures calibrated margins and realistic dependence reconstruction. We evaluate COBASE on multi-site 2-meter temperature forecasts from the ALADIN-LAEF ensemble over Austria and on joint forecasts of temperature and dew point temperature from the ECMWF system in the Netherlands. Across all regions, COBASE variants consistently outperform traditional copula-based approaches, such as GCA, and achieve performance on par with state-of-the-art nonparametric methods like SimSchaake and ECC, with only minimal differences across settings. These results position COBASE as a competitive and robust alternative for multivariate ensemble postprocessing, offering a principled bridge between parametric and nonparametric dependence reconstruction.
keywords:
Statistical post-processing of weather forecasts , Two-step parametric correction , Dependence modeling and copulas , ALADIN-LAEF ensemble system , ECMWF.
journal: arXiv
1 Introduction
The forecasting skills of the physical numerical weather prediction (NWP) models have been consistently improved in recent years (e.g., Bauer et al., 2015). Efforts in research have focused on enhancing physical parameterizations, assimilating non-conventional data, and boosting spatial resolution. Despite the progress made, deterministic and probabilistic NWP forecasts still show systematic errors and underdispersion. Potential causes of errors are imperfect measurements (which also serve as initial conditions, e.g., Cardinali, 2009), imprecise parametrization of physical equations (e.g., Bechtold et al., 2014) and inherent unpredictability due to chaotic dynamics in atmospheric flows (e.g., Lorenz, 1963). To account for such model errors, calibration approaches based on statistical or machine learning post-processing techniques have been developed in the last years: see, for example, Gneiting et al. (2005); Raftery et al. (2005); Möller et al. (2013); Schefzik et al. (2013); Taillardat et al. (2016); Schefzik (2016); Scheuerer and Hamill (2018); Rasp and Lerch (2018); Lerch et al. (2020); Vannitsem et al. (2021). The standard statistical methods for the post-processing of raw ensemble forecasts are Ensemble Model Output Statistics (EMOS) (Gneiting et al., 2005) and Bayesian Model Averaging (BMA) (Raftery et al., 2005). These techniques work well in practice, but are designed to improve predictions for only one variable at a time, e.g., for a fixed weather variable at a single location and specific lead-time. As a result, such methods fail to maintain the inter-variable and spatio-temporal dependencies of the forecasts (e.g., Schefzik, 2016). This poses challenges for forecasting high-impact weather events such as floods (Schaake et al., 2010), heat waves (e.g., Baran et al., 2020), or the onset of hurricanes, where numerous weather factors intricately interact. To address this problem, multivariate postprocessing methods have been developed. The most popular techniques are based on empirical copulas (e.g., Schefzik et al., 2013; Lakatos et al., 2023). Namely, these are two-step approaches that use standard EMOS or BMA to correct the univariate forecast and a reshuffling technique to reconstruct the lost dependencies in the samples that are drawn from the univariate distributions. There are several ways to sample from the post-processed distribution, including drawing random samples, uniform quantiles, or stratified samples. Uniform quantiles or stratified samples generally perform better than random samples, particularly for extremes (Wilks, 2015; Hu et al., 2016). Examples of reshuffling methods are Schaake Shuffle (Clark et al., 2004), Ensemble Copula Coupling (ECC) (Schefzik et al., 2013), and Sim Schaake (Schefzik, 2016). The common ground of these methods is that the dependence structure can be obtained from a reference sample, which can be based on past observations (for Schaake Shuffle and Sim Schaake) or on the raw forecasts (for ECC). However, these methods have some important limitations as they require a large number of past observations or a reliable, exchangeable ensemble system that simulates a realistic dependence structure to be effective. These conditions are not always satisfied as is the case for the ALADIN-LAEF system used in Austria (Wang et al., 2011). In such cases, multivariate parametric approaches that do not require extensive data archives nor do not assume exchangeability, could be better suited for multivariate post-processing.
Multivariate parametric approaches for the statistical postprocessing of weather forecasts already exist (e.g., Möller et al., 2013; Schefzik and Möller, 2018; Jobst et al., 2024; Lerch et al., 2020). Parametric methods with advanced copula models improve the dependence structure of the corrected sample but often fail to maintain a suitable marginal calibration as they can only be used with random sampling which might perform poorly compared to uniform quantile sampling (Wilks, 2015). This is a problem that puts copula-based methods at a disadvantage when compared to non-parametric two-step approaches which can be applied to any univariate corrected sample. We believe that this limitation has hindered the potential of multivariate parametric postprocessing in both academic research and operational use.
This paper introduces a novel approach aimed at fully harnessing the potential of two-step parametric correction techniques. Specifically, we introduce a new copula-based shuffling technique (COBASE) that leverages the adaptability of parametric copulas while preserving the marginal calibration. We present the methodology through two case studies, similar to those examined in Perrone et al. (2020) and Whan et al. (2021). Our attention is directed towards spatial dependencies in 2-meter temperature predictions using the ALADIN-LAEF system in Austria and inter-variable and spatial dependencies in 2-meter temperature and dew-point temperature predictions the Netherlands. The case studies demonstrate the applicability of our method in different climates, with different numerical weather prediction systems (regional and global) and number of dimensions. Our approach significantly outperforms the state-of-the-art multivariate parametric methods, and it provides a viable way to enhance the performance of any parametric copula-based method in the post-processing context.
The paper is organized as follows. In Section 2, we present the two numerical weather prediction systems and our case study settings. In Section 3, we review existing non-parametric and parametric methods for post-processing, discuss the limitations of current approaches, and introduce COBASE, our new methodology. Section 4 tests COBASE in our case studies. Finally, we draw conclusions and end with a general discussion in Section 5.
2 Data
2.1 Case study 1: spatial dependencies in Austria
First, we consider a case study based on the ALADIN-LAEF model (Wang et al., 2011), which is the predecessor of the currently operational Austrian ensemble forecasting system C-LAEF (Wastl et al., 2021). The ALADIN-LAEF system has been optimized for the Austrian atmsopheric conditions and was operational from 2013 to 2022. The migration to the successor model C-LAEF was carried out in parallel in 2022 and put into operation in 2023. Although C-LAEF replaced ALADIN-LAEF operationally in 2023, historical raw forecasts from C-LAEF remain scarce. Therefore, ALADIN-LAEF data were used in this study.
The ALADIN-LAEF ensemble included a total of 17 ensemble members, comprising 16 perturbed members that are guided by 16 ECMWF-EPS members, along with one control run. ECMWF members acted as boundary conditions for ALADIN-LAEF, which means that each ALADIN-LAEF member was based on an ECMWF member. However, the physical parameterization within the ALADIN-LAEF ensemble varied, and the initial observations were perturbed in distinct ways. As a result, ALADIN-LAEF was not an exchangeable system and did not contain any of the original and unmodified members of the ECMWF. ALADIN-LAEF was set to run with a horizontal resolution of 10.9 km using a Lambert conformal projection. For the vertical structure, it incorporated 45 hybrid pressure-coordinate levels aligned with the terrain, averaging nine levels in the lowest 1000 km from the surface. The system underwent initialization twice a day at 0000 and 1200 UTC, offering hourly lead-times extending up to 72 hours into the future. The ALADIN-LAEF forecast data was subjected to bilinear interpolation to align with the locations of the TAWES (TeilAutomatische WEtterStation) network in Austria. The TAWES network, utilized for weather observations across Austria, consists of approximately 300 sites that cover almost every altitude and climatic region of the nation. This research examines 2-meter temperature (T2m) data collected from specific TAWES locations. Here, we analyze the same setup and dataset used in Perrone et al. (2020). Specifically, we examine three groups of stations over a time span extending from January 2014 to May 2018. Each group includes three sites, demonstrating a unique feature of interest as detailed below.
  1. 1.
    LAEF - Single Valley. The initial group comprises three stations situated in proximity within a valley. The stations chosen are Sonnblick (11343), Kolm Saigurn (11344), and Rauris (11346).
  2. 2.
    LAEF - Mountain Peaks. The second group includes the stations located on the tops of three mountains. We focus on Patscherkofel (11126), Pitzal (11316), and Sonnblick (11343).
  3. 3.
    LAEF - Randomly Chosen. The third group comprises three stations that are each randomly picked from separate regions within Austria. These are Wien Hohe Warte (11035), Innsbruck (11320), and Graz (11290).
In Single Valley, we note that the stations are located near each other. In contrast, the stations within the Mountain Peaks and Randomly Chosen categories are distributed across regions spanning hundreds of kilometers. This research aims to assess spatial dependencies while assuming that prediction horizons are fixed. In particular, the study is based on a fixed lead-time of 24h.
2.2 Case study 2: inter-variable and spatial dependencies in the Netherlands
For this case study, we take operational forecasts of 2-meter temperature (T2m) and dew-point temperature (DPT) from the global ensemble prediction system run by the European Center for Medium-Range Weather Forecasting (ECMWF). We use the 51-members that are initialized each day at 12 UTC during the period 2021-2024, and take the forecasts that are valid at +24 hours. There were several model upgrades during this time, including a change in horizontal resolution from  18km to  9km in 2023. The use of a rolling training window, as explained in Section 3, reduces the impact of these model changes.
We take observations of T2m and DPT from the KNMI network of automatic weather stations in the Netherlands, similarly to Whan et al. (2021). We match the forecasts to the observations with cubic interpolation. For our analysis, we consider one group as detailed below.
  1. 4.
    ECMWF ENS. The last group comprises six stations from the total set of stations in the Netherlands chosen to be evenly distributed across the country. Namely, we selected De Kooy (235), Schiphol (240), De Bilt (260), Groningen (280), Vlissingen (310), and Maastricht (380).
3 Methods
This section outlines the two-step approaches currently considered state-of-the-art in multivariate statistical post-processing. The key idea of these methods is to separate the correction of forecasts into two parts, one that addresses the marginal forecasts, and another that pertains the dependence structure. In practical terms, this involves initially applying univariate post-processing techniques to each margin separately, meaning that we address each weather variable, lead time, and station on an individual basis. Afterwards, we restore the missing dependencies by applying the theory of copulas.
Copulas are cumulative joint distributions with uniform margins on $[0,1]$. The theoretical foundation of copula theory is Sklar’s theorem (Sklar, 1959), which states that the $d$-dimensional multivariate cumulative distribution function $F$ of a random vector $(X_{1},\ldots,X_{d})$ with marginals $F_{1},\ldots,F_{d}$ can be expressed as the composition of a copula $C$ and univariate marginal distributions, i.e., $F(x_{1},\ldots,x_{d})=C(F_{1}(x_{1}),\ldots,F_{d}(x_{d}))$. Therefore, copulas can identify the dependence structure of a phenomenon without considering its marginal distributions. In the post-processing context, they can be used in the second step to reintroduce a dependence structure which gets lost in the univariate post-processing and sampling step. In the remainder of the section, we first introduce the univariate correction method used in our study, followed by the nonparametric and parametric copula-based approaches, and our newly proposed method COBASE. A summary of the specifications of the methods used with the corresponding labels is given in Table 2.
3.1 Univariate Post-processing: Ensemble Model Output Statistics (EMOS)
One widely adopted method for univariate correction is Ensemble Model Output Statistics (EMOS) (Gneiting et al., 2005). EMOS employs ensemble forecasts as covariates to adjust the parameters of a Gaussian distribution, aiming to correct errors in the mean as well as in the forecast uncertainty. In this paper, we follow the same approach as in Perrone et al. (2020) and define the regression model as:
$y$ $\sim\mathcal{N}(\mu,\sigma^{2})$ (1)
$\mu$ $=\alpha_{0}+\alpha_{1}\cdot m$ (2)
$\sigma^{2}$ $=\beta_{0}+\beta_{1}\cdot s^{2}$ (3)
[…]
$\text{CRPS}(F_{ens},y)=\dfrac{1}{M}\sumop\displaylimits\limits_{m=1}^{M}|x_{m}-y|-\dfrac{1}{2M^{2}}\sumop\displaylimits\limits_{n=1}^{M}\sumop\displaylimits\limits_{m=1}^{M}|x_{n}-x_{m}|.$ (4)
While EMOS provides effective univariate corrections, it does not capture dependencies across multiple variables or locations, motivating the use of multivariate post-processing methods. In the following section, we review non-parametric empirical copula-based postprocessing approaches.
3.2 Nonparametric multivariate post-processing: Empirical-copula methods
This section reviews the most common two-step nonparametric techniques for statistical post-processing. All of these methods are based on empirical copulas, which offer an effective way to incorporate the dependence structure into weather forecasting problems without imposing any parametric assumptions (Schefzik, 2015). An empirical copula is essentially a discrete form of a copula $C$, known as a discrete copula, derived from a given multivariate sample (Schefzik, 2015). As such, it encodes the dependence structure of the reference sample through the corresponding ranks. In weather forecasting, empirical copulas can be used to generate adjusted multivariate forecasts that integrate dependencies taken from a reference rank structure. This idea was first introduced in the Schaake Shuffle method (Clark et al., 2004), which uses the rank structure obtained from historical observations. Then, this concept was refined and new methods were developed, the most popular being ECC (Schefzik et al., 2013) and Sim Schaake (Schefzik, 2016). ECC reconstructs the dependence according to the raw ensemble forecasts. Sim Schaake combines the ideas of ECC and Schaake shuffle: First, it selects past days where the ensemble forecasts are similar to the verification day and then uses the selected observations to reconstruct the dependence.
[…]
where $\bar{x}^{l,\tau}$ and $s_{l,\tau}$ are the empirical mean and standard deviation of the ensemble forecast $\mathbf{x}^{\tau}$ for date $\tau$, i.e., they are defined as follows:
$\bar{x}^{l,\tau}:=\dfrac{1}{M}\sumop\displaylimits_{m=1}^{M}x_{m}^{l,\tau},\quad s_{l,\tau}:=\sqrt{\dfrac{1}{M-1}\sumop\displaylimits_{m=1}^{M}\left(x_{m}^{l,\tau}-\bar{x}^{l,\tau}\right)^{2}}.$
In addition to the construction of the reference dependence structure $DC_{ref}$, another point that can generally vary is the sampling technique used in S3 of Algorithm 1. In Schefzik et al. (2013), various sampling techniques have been discussed. One could simply take equidistant quantiles of step $N+1$ from the univariate corrected distribution $F_{l}$:
$\tilde{x}^{l}_{1}=F^{-1}_{l}\left(\dfrac{1}{N+1}\right),\;\ldots,\;\tilde{x}^{l}_{N}=F^{-1}_{l}\left(\dfrac{N}{N+1}\right).$ (6)
There is a general consensus in the literature that the most natural way to construct a discrete sample of size $N$ by uniform quantiles is also the most effective in this context (e.g. Wilks, 2015). This is true especially when uniform quantiles are compared with randomly drawn samples. Due to this and in line with Perrone et al. (2020), we use uniform quantiles in our study.
Algorithm 1 Two-step rank-based postprocessing methods
S1: Derive a reference dependence structure $DC_{ref}$;
S2: Apply EMOS and obtain a corrected distribution $F_{l}$ for each margin $l=1,\ldots,d$;
[…]
S4: Shuffle the corrected univariate samples according to the ranks of $DC_{ref}$.
In our case studies, we ensure a fair comparison with ECC by setting the size of the ensemble at $N=17$ for the ALADIN-LAEF groups and $N=51$ for the ECMWF data. For the Sim Schaake and Schaake Shuffle methods, we use all available observations and raw forecasts, including past and future dates (approximately four years), to ensure a consistent number of reference dates throughout the test set. Specifically for the Schaake Shuffle, we select observations randomly from a 14-day window before and after the forecast date to preserve the seasonal climatology.
The main properties of the empirical copula-based approaches described here are summarized in Table 1. As emphasized in the table, the main limitations of these approaches are as follows.
  • (a)
    The necessity for extensive historical data to more accurately illustrate the dependence influences both the Schaake Shuffle and Sim Schaake.
  • (b)
    The requirement for extensive historical raw forecasts is crucial for the similarity criterion, impacting Sim Schaake.
  • (c)
    The requirement that the sample size of the corrected forecasts ($N$) must equal the number of ensemble members ($M$), imposes a significant limitation on the applicability of ECC.
Although these limitations might seem less relevant in an operational context, they may have a significant impact in practice. An example is Case Study 1 in Austria, where the post-processing of the operational system C-LAEF could not be performed due to the lack of historical raw forecasts and historical observations. Furthermore, for ALADIN-LAEF, the limited ensemble size of only $17$ members posed a substantial constraint on the application of the ECC method. Consequently, this case study demonstrates a scenario in which the shortcomings of these methods might justify a transition to alternative approaches, such as parametric multivariate postprocessing approaches. We introduce these approaches in the following section after providing an overview of how multivariate forecasts are evaluated in this study.
[…]
$\text{ES}=\dfrac{1}{M}\sumop\displaylimits\limits_{m=1}^{M}\|x_{m}-y\|-\dfrac{1}{2M^{2}}\sumop\displaylimits\limits_{\nu=1}^{M}\sumop\displaylimits\limits_{m=1}^{M}\|x_{\nu}-x_{m}\|$ (7)
In addition, the Variogram Score (VS) (Scheuerer and Hamill, 2015) explicitly addresses miscalibrations in the correlation structure of multivariate forecasts. In the ensemble setting, it is defined as
$VS=\sumop\displaylimits_{i=1}^{d}\sumop\displaylimits_{j=1}^{d}{w}_{i,j}\left(\left|{y}_{i}-{y}_{j}\right|^{p}-\frac{1}{m}\sumop\displaylimits_{k=1}^{m}\left|x^{(k)}_{i}-x^{(k)}_{j}\right|^{p}\right)^{2}.$ (8)
where $w_{i,j}$ are weights and $p>0$ is the order of the score. Because the VS compares pairwise differences between variables in the observation and forecast, it is particularly sensitive to errors in the dependence structure. In our case study, we use both the Energy Score and the Variogram Score to assess and compare the performance of the different multivariate post-processing methods. These metrics provide complementary information, evaluating both the overall accuracy of the multivariate forecasts and their ability to represent dependence structures. We used the variant with no weights and $p=1$ for the Variogram Scores, but we obtained consistent results for other variants as well.
3.3 Existing parametric multivariate post-processing
The reordering strategies discussed so far are, in fact, nonparametric. Specifically, within the context of Sklar’s theorem, the copula $C$ is represented by the empirical copula $DC_{ref}$ derived from a reference dependence structure that varies for each method. Another way to proceed would be to follow a two-step parametric approach. Namely, we fit a parametric copula $C$ to the training data and use it to generate a corrected multivariate sample with prescribed marginal distributions obtained through EMOS.
[…]
Hence, the observations are transformed into normally distributed variables that are used to fit the covariance matrix. This covariance matrix can be thought of as the copula parameter that models all dependencies and is used to create the forecasts. Therefore, GCA can be considered as a genuine copula-based approach where $C$ is the Gaussian copula.
The copula-based parametric correction steps for post-processing are summarized in Algorithm 2, which illustrates that any copula model can be used in place of the Gaussian copula. In our case studies, we estimate the copula parameters for standard parametric copula models using a 30-day rolling window, consistent with the EMOS procedure.
We notice that there has been limited research on the application of alternative copulas in the post-processing context, with notable exceptions such as Jobst et al. (2024) and Flos (2022). As highlighted in Table 1, while parametric copula-based methods exhibit considerable flexibility, a key limitation lies in their reliance on random sampling. Unlike structured techniques, such as the use of uniform quantiles, random draws provide a sparse and noisy representation of the univariate distributions. Since copula methods always generate corrected samples through random draws from a multivariate distribution, the resulting margins inherit this inefficiency. Consequently, univariate skill often suffers even when multivariate dependencies are well captured. In the next section, we provide evidence that this issue is substantial and would affect the performance of any multivariate parametric post-processing method.
Algorithm 2 Two-step copula-based postprocessing methods
S1: Estimate the parameters $\alpha$ of a copula model $C_{\alpha}$;
S2: Apply EMOS and obtain a corrected distribution $F_{l}$ for each margin $l=1,\ldots,d$;
[…]
Figure 1: Boxplots of Diebold-Mariano test statistics for CRPS of T2m, comparing EMOS-Q against EMOS-R (baseline). Results are presented for the Austrian and Dutch stations used in this study.
Figure 2: Boxplots of Diebold-Mariano test statistics to compare ES and VS of Sim Schaake with uniform quantile sampling (SimSSh) versus Sim Schaake with randomly drawn margins (SimSSh-R, baseline). Results are presented for T2m for the three groups of Austrian stations and for T2m and DPT for the group of Dutch stations.
3.3.1 Limitations of univariate random sampling in statistical post-processing
As mentioned earlier, the reliance of some univariate and multivariate post-processing techniques on random sampling can lead to poorly calibrated margins. To illustrate this, we compare the CRPS of T2m obtained for all stations in the study (Figure 1) using EMOS-R, i.e., random draws from the corrected Gaussian distribution, as the baseline. Figure 1 shows that drawing uniform quantiles instead (EMOS-Q) consistently outperforms EMOS-R across all stations.
We further examine how univariate random sampling affects multivariate corrections by comparing the Energy Score and the Variogram Score for the Sim Schaake method with different marginal samples, i.e., EMOS-R versus EMOS-Q. Essentially, the two methods compared in Figure 2 share the same dependence structure and only differ in the marginal sampling technique. The results show that Sim Schaake with randomly drawn margins (SimSSh-R) consistently performs worse than Sim Schaake with uniform quantile margins (SimSSh) in both scores across all groups, demonstrating that poor univariate sampling propagates into multivariate performance.
This investigation highlights how random sampling of marginal distributions can negatively affect multivariate corrections. Since copula-based techniques fundamentally rely on random draws from parametric multivariate distributions, this limitation may systematically compromise their performance. To overcome this limitation, we develop a new method, COBASE, which is presented in the following section.
3.4 A new copula-based shuffling method for ensemble postprocessing: COBASE
The proposed Copula-Based Shuffling of Ensemble (COBASE) method extends existing copula-based postprocessing techniques by introducing a shuffling of corrected univariate forecasts that replaces random marginal sampling. This approach preserves both realistic marginal behavior and the dependence structure defined by the chosen copula. The procedure is summarized in Algorithm 3. First, a parametric copula $C_{\alpha}$ is fitted to the data to capture the underlying dependence structure (Step S1). A random sample is then drawn from this copula to serve as the reference dependence structure $C_{ref}$ (Step S2). Next, standard univariate postprocessing, such as EMOS, is applied to each margin $l=1,\ldots,d$ to obtain corrected distributions $F_{l}$ (Step S3). Independent samples of size $N$ are drawn from these corrected margins using uniform quantiles (Step S4). This ensures that the marginal distributions are well-calibrated, including at the extremes, avoiding the shortcomings of random draws from a parametric multivariate distribution. Importantly, this step is flexible: any univariate sampling technique can be used, not necessarily uniform quantiles, for example, stratified sampling, allowing additional control over the sampled distributions. Finally, the univariate samples are shuffled according to the rank structure of $C_{ref}$ (Step S5). By doing so, COBASE preserves the dependence structure depicted by the parametric copula while leveraging the improved marginal calibration of advanced univariate sampling through, for example, uniform quantiles. As previously discussed, Table 1 provides an overview of the main multivariate post-processing methods and their key properties. In the table, COBASE stands out by combining the strengths of parametric and non-parametric two-step approaches: The parametric copula captures dependence, while uniform quantile sampling ensures well-calibrated margins. In the next section, we evaluate COBASE in two real-world case studies, comparing it with state-of-the-art methods to highlight its practical advantages. In the two case studies, we estimate the copula parameters for COBASE using a 30-day rolling window, consistent with the EMOS procedure and the standard copula models approaches.
Algorithm 3 Our proposed approach - COBASE
S1: Estimate the parameters $\alpha$ of a copula model $C_{\alpha}$;
[…]
S3: Apply EMOS and obtain a corrected distribution $F_{l}$ for each margin $l=1,\ldots,d$;
[…]
S5: Shuffle the corrected univariate samples according to the ranks of $C_{ref}$.
Method Approach Dependence Free from long historical data Free from historical raw forecasts Flexible marginal sampling No restriction on the sample Size
Sim Schaake Shuffling Empirical copula (from past obs) × ×
Schaake shuffle Shuffling Empirical copula (from past obs) ×
ECC Shuffling Empirical copula (from raw forecasts) ×
Copula methods (parametric) Parametric fit (copula model) Parametric copula model (rolling window) ×
COBASE Parametric fit (copula model) + shuffling Parametric copula model (rolling window)
Table 1: Summary of the multivariate post-processing methods analyzed in this work, alongside their main characteristics.
Figure 3: Boxplots of Diebold-Mariano test statistics for the case studies. Here, we compare ES and VS of Schaake Shuffle (SSh), Sim Schaake (SimSSh), ECC, and GCA with COBASE-GCA as the baseline. Values above the gray band indicate better performances compared to the baseline. Results are presented for T2m for the three groups of Austrian stations and for T2m and DPT for the group of Dutch stations.
4 Results
We evaluate the proposed COBASE method in comparison with state-of-the-art post-processing approaches on the two real-world case studies described in Section 2. Although the focus of this work is not on univariate correction, CRPS results for 2-meter temperature across all stations are included in Table 3 for completeness. These results show that EMOS-Q, which uses uniform quantile sampling, consistently outperforms methods based on random sampling, clearly indicating that quantile-based approaches are generally preferable. The main focus of this section is on multivariate forecast performance, assessed using the Energy Score (ES) and the Variogram Score (VS). We begin by evaluating COBASE-GCA, the variant of COBASE using a Gaussian copula. COBASE-GCA is compared with Sim Schaake, Schaake Shuffle, ECC, and standard GCA.
The results, presented as boxplots of Diebold–Mariano (DM) test statistics with COBASE-GCA as reference, are shown in Figure 3, with the corresponding ES and VS values reported in Table 4. These findings show that COBASE-GCA consistently outperforms standard GCA for both ES and VS for all groups of stations. Additionally, COBASE-GCA generally exhibits comparable performance across the different groups of stations compared to the other empirical copula-based methods. In particular, COBASE-GCA outperforms all the other methods in LAEF (Mountain Peaks). COBASE-GCA outperforms ECC and Schaake shuffle in LAEF (Single Valley) and LAEF (Randomly chosen), with Sim Schaake being preferable in those instances. Finally, for ECMWF ENS data, COBASE-GCA achieves the best ES, while ECC performs best in terms of VS.
We also evaluate COBASE with alternative copula families, namely, Clayton, Frank, and Gumbel, to assess its use and evaluate performance under different (non-Gaussian) dependence structures. The corresponding DM test results are displayed in Figure 4. These results indicate that the COBASE variants consistently outperform their corresponding standard copula-based implementation. This conclusion is further supported by the ES and VS scores, also reported in Table 4.
Figure 4: Boxplots of Diebold-Mariano test statistics for the case studies. Here we compare ES and VS of four copula models (baseline) versus their COBASE variants. Values above the gray band indicate better performances compared to the baseline.
5 Discussion and conclusion
This paper introduces COBASE, a new multivariate post-processing method that combines the structure of parametric copula models with the flexibility of non-parametric methods. By decoupling the sampling of margins from the copula-based dependence structure, COBASE addresses a key issue in the copula-based approaches for post-processing, namely, the degradation in forecast quality caused by random sampling from marginal distributions.
A key strength of COBASE is its flexibility. The method is compatible with any parametric copula family, allowing it to adapt to a wide range of dependence structures. In addition, it does not impose constraints on how univariate margins are sampled. While uniform quantiles are used here, the method is fully compatible with alternative sampling strategies, such as stratified sampling, making it adaptable to different forecasting needs.
Our results from two real-world case studies in Austria and the Netherlands show that COBASE consistently improves upon standard copula-based methods, including GCA, and outperforms Schaake Shuffle. Moreover, COBASE achieves performance comparable to widely used non-parametric techniques such as Sim Schaake and ECC, while avoiding their structural limitations and extensive data requirements.
A closer look at the scores in Table 4 reveals that COBASE-GCA performs similarly to Sim Schaake, even in the station groups where the latter shows a slight advantage. It is important to note, however, that the comparison is not entirely fair: Sim Schaake and Schaake Shuffle are trained on four years of data, whereas COBASE, like EMOS, uses only a 30-day rolling training window. This shorter training horizon makes COBASE substantially more flexible and easier to apply in operational contexts. For example, COBASE would be applicable to C-LAEF, while Sim Schaake would not be suitable in that case due to the lack of past raw forecasts. Additionally, COBASE shows robust performance even in relatively high-dimensional settings, such as the ECMWF ENS group with 12 variables, and no dimension-specific tuning.
The added value of COBASE becomes particularly evident in comparison to that of its corresponding standard copula-based variant. In this case, the performance gap is much more pronounced, confirming that the observed improvements stem from the structure of the COBASE method itself rather than from the specific copula family employed. This supports the main goal of the study: Not to identify the single best-performing copula, but to demonstrate that, across different configurations, COBASE consistently enhances the effectiveness of copula-based post-processing. In this sense, COBASE represents a more efficient, flexible, and broadly applicable alternative to conventional parametric multivariate approaches.
In the future, we plan to further investigate several aspects that could enhance the performance and versatility of the method. First, the quality of the estimated dependence structure depends strongly on both the choice of copula family and the accuracy of parameter estimation, which can be affected by limited sample sizes. In this study, we focused on temperature, a variable that is approximately Gaussian, so the Gaussian copula (GCA) already provided a good representation of dependence. For other variables, such as wind speed or precipitation, where deviations from Gaussianity are more pronounced or extremes are of greater interest, exploring alternative copula families or model selection strategies may yield further improvements.
Second, future work could aim to make COBASE more adaptive by allowing the dependence structure, and thus the copula family, to vary dynamically with the forecast date or prevailing meteorological regime. Such adaptivity could enable the method to better capture time-varying or state-dependent relationships between variables. However, implementing dynamically varying copulas would also increase computational demands, so developing efficient strategies for adaptive copula selection or updating will be essential for practical applications.
Finally, while COBASE avoids the limitations of random marginal sampling, it still relies on random draws from the fitted copula to represent the multivariate structure. Future research could explore alternative multivariate sampling strategies to reduce sampling variability and better capture the characteristics of the dependence structure, especially in the tails of the multivariate distributions. These avenues offer promising directions for further strengthening COBASE’s performance across a wider range of postprocessing tasks.
To conclude, taken together, our findings demonstrate that COBASE already provides a simple, flexible, and effective framework for multivariate ensemble post-processing. By combining the strengths of parametric and non-parametric approaches, it improves forecast skill without requiring long historical records or raw ensemble data. The COBASE methodology is also particularly well suited for research on post-processing of multivariate dependence, as it allows a clear separation between the evaluation of the dependence structure and the marginals, thereby facilitating a better identification of potential improvements in the dependence component itself. These properties make COBASE particularly well-suited for operational settings with limited data or for newly developed forecasting systems, such as the C-LAEF system in Austria.
Building on this foundation, future work will focus on adjusting the method to higher-dimensional settings, improving copula selection strategies, and enhancing the multivariate sampling step. With these developments, COBASE has the potential to become a robust and scalable solution for a wide range of ensemble post-processing applications.
Data availability statement
The datasets analyzed during this study are subject to restrictions and are therefore not publicly available. To facilitate reproducibility, all analysis scripts and mock data that closely replicate the original datasets, allowing reproduction of the figures and tables, are available at https://github.com/elisaperrone/COBASE_github.
Appendix A Abbreviations
Method Label
Univariate postprocessing
EMOS (random sampling) EMOS-R
EMOS (uniform quantile sampling) EMOS-Q
Multivariate postprocessing
Shuffling:
Sim Schaake + EMOS-R SimSSh-R
Sim Schaake + EMOS-Q SimSSh
Schaake Shuffle + EMOS-Q SSh
ECC + EMOS-Q ECC
Copula approaches (parametric):
Gaussian Copula GCA
Clayton Copula Clayton
Frank Copula Frank
Gumbel Copula Gumbel
Copula + Shuffling:
COBASE - Gaussian Copula + EMOS-Q COBASE-GCA
COBASE - Clayton Copula + EMOS-Q COBASE-Clayton
COBASE - Frank Copula + EMOS-Q COBASE-Frank
COBASE - Gumbel Copula + EMOS-Q COBASE-Gumbel
Table 2: Overview of the statistical post-processing methods evaluated in the study, grouped by approach type, with their corresponding labels.
Appendix B Scores
B.1 CRPS
Station EMOS-R EMOS-Q GCA Gumbel Clayton Frank
Wien Hohe Warte (11035) 0.8376 0.8055 0.8417 0.8457 0.8409 0.8548
Patscherkofel (11126) 1.3846 1.3429 1.3920 1.3903 1.3939 1.3883
Graz (11290) 0.9166 0.8848 0.9210 0.9244 0.9195 0.9260
Pitzal (11316) 1.0912 1.0460 1.0831 1.0791 1.0742 1.0838
Innsbruck (11320) 1.2000 1.1515 1.1922 1.1871 1.1946 1.2020
Sonnblick (11343) 0.9650 0.9257 0.9624 0.9640 0.9530 0.9522
Kolm Saigurn (11344) 1.3982 1.3445 1.4146 1.3989 1.3920 1.3938
Rauris (11346) 1.5313 1.4726 1.5289 1.5421 1.5542 1.5514
De Kooy (235) 0.6071 0.6030 0.6113 0.6106 0.6133 0.6096
Schiphol (240) 0.5901 0.5873 0.5957 0.5956 0.5983 0.5973
De Bilt (260) 0.5772 0.5695 0.5777 0.5786 0.5748 0.5758
Groningen (280) 0.6068 0.5939 0.6026 0.6045 0.6029 0.5987
Vlissingen (310) 0.6521 0.6421 0.6544 0.6500 0.6512 0.6522
Maastricht (380) 0.6325 0.6232 0.6316 0.6310 0.6293 0.6335
Table 3: CRPS values for all individual stations and postprocessing methods for 2-meter temperature. To avoid redundancy, results are presented for EMOS-R, EMOS-Q, GCA, Gumbel, Clayton and Frank methods only. All the other approaches considered in this study rely on reshuffling the marginals of these methods and produce identical marginal performance. The gray cells indicate the best scores.
B.2 ES and VS
Method Regions and scores
LAEF (Single Valley) LAEF (Mountain Peaks) LAEF (Randomly Chosen) ECMWF ENS
ES VS ES VS ES VS ES VS
Raw Ensemble 7.6462 239.5952 4.3578 57.2592 7.3904 202.6426 3.1251 354.8837
SSh 2.6007 50.9879 2.2322 20.0152 1.9618 19.7619 2.7629 303.3426
SimSSh-R 2.6187 50.1894 2.2837 20.1669 1.9824 19.6075 2.7084 305.0797
SimSSh 2.5292 47.8356 2.2079 19.3942 1.9325 18.6810 2.6923 299.1751
ECC 2.6575 53.2521 2.2130 19.5822 1.9513 19.3800 2.6989 298.7708
GCA 2.6180 50.1523 2.2569 19.5415 1.9860 19.8129 2.7078 303.4243
Clayton 2.6265 50.4588 2.2530 19.6612 1.9862 19.4570 2.7137 303.2181
Frank 2.6254 50.4119 2.2593 19.7900 2.0008 19.7574 2.7093 304.4875
Gumbel 2.6363 50.6686 2.2587 20.0026 1.9843 19.5161 2.7089 303.1059
COBASE-GCA 2.5344 48.3626 2.2073 19.0118 1.9363 18.8943 2.6908 299.2300
COBASE-Clayton 2.5501 48.9828 2.2080 19.1085 1.9355 18.7195 2.6921 299.4446
COBASE-Frank 2.5504 48.9279 2.2056 19.1347 1.9344 18.7868 2.6905 299.7904
COBASE-Gumbel 2.5601 49.4700 2.2080 19.1400 1.9345 18.7907 2.6926 299.3518
Table 4: ES and VS for T2m (LAEF) and T2m and DPT (ECMWF ENS) for the different multivariate postprocessing methods. The gray cells indicate the best scores.

### v2

@@
 keywords:
-Statistical post-processing of weather forecasts , Two-step parametric correction , Dependence modeling and copulas , ALADIN-LAEF ensemble system , ECMWF.
+Statistical postprocessing of weather forecasts , Two-step parametric correction , Dependence modeling and copulas , ALADIN-LAEF ensemble system , ECMWF.
 journal: arXiv
@@
 Potential causes of errors are imperfect measurements (which also serve as initial conditions, e.g., Cardinali, 2009), imprecise parametrization of physical equations (e.g., Bechtold et al., 2014) and inherent unpredictability due to chaotic dynamics in atmospheric flows (e.g., Lorenz, 1963).
-To account for such model errors, calibration approaches based on statistical or machine learning post-processing techniques have been developed in the last years: see, for example, Gneiting et al.
+To account for such model errors, calibration approaches based on statistical or machine learning postprocessing techniques have been developed in the last years: see, for example, Gneiting et al.
 (2005); Raftery et al.
@@
 (2021).
-The standard statistical methods for the post-processing of raw ensemble forecasts are Ensemble Model Output Statistics (EMOS) (Gneiting et al., 2005) and Bayesian Model Averaging (BMA) (Raftery et al., 2005).
+The standard statistical methods for the postprocessing of raw ensemble forecasts are Ensemble Model Output Statistics (EMOS) (Gneiting et al., 2005) and Bayesian Model Averaging (BMA) (Raftery et al., 2005).
 These techniques work well in practice, but are designed to improve predictions for only one variable at a time, e.g., for a fixed weather variable at a single location and specific lead-time.
@@
 The common ground of these methods is that the dependence structure can be obtained from a reference sample, which can be based on past observations (for Schaake Shuffle and Sim Schaake) or on the raw forecasts (for ECC).
-However, these methods have some important limitations as they require a large number of past observations or a reliable, exchangeable ensemble system that simulates a realistic dependence structure to be effective.
+However, these methods have some important limitations as they require on sufficiently representative past observations or a reliable, exchangeable ensemble system that simulates a realistic dependence structure to be effective.
 These conditions are not always satisfied as is the case for the ALADIN-LAEF system used in Austria (Wang et al., 2011).
-In such cases, multivariate parametric approaches that do not require extensive data archives nor do not assume exchangeability, could be better suited for multivariate post-processing.
+In such cases, multivariate parametric approaches that do not require extensive data archives nor do not assume exchangeability, could be better suited for multivariate postprocessing.
 Multivariate parametric approaches for the statistical postprocessing of weather forecasts already exist (e.g., Möller et al., 2013; Schefzik and Möller, 2018; Jobst et al., 2024; Lerch et al., 2020).
 Parametric methods with advanced copula models improve the dependence structure of the corrected sample but often fail to maintain a suitable marginal calibration as they can only be used with random sampling which might perform poorly compared to uniform quantile sampling (Wilks, 2015).
-This is a problem that puts copula-based methods at a disadvantage when compared to non-parametric two-step approaches which can be applied to any univariate corrected sample.
+This is a problem that puts copula-based methods at a disadvantage when compared to nonparametric two-step approaches which can be applied to any univariate corrected sample.
 We believe that this limitation has hindered the potential of multivariate parametric postprocessing in both academic research and operational use.
@@
 The case studies demonstrate the applicability of our method in different climates, with different numerical weather prediction systems (regional and global) and number of dimensions.
-Our approach significantly outperforms the state-of-the-art multivariate parametric methods, and it provides a viable way to enhance the performance of any parametric copula-based method in the post-processing context.
+Our approach significantly outperforms the state-of-the-art multivariate parametric methods, and it provides a viable way to enhance the performance of any parametric copula-based method in the postprocessing context.
 The paper is organized as follows.
 In Section 2, we present the two numerical weather prediction systems and our case study settings.
-In Section 3, we review existing non-parametric and parametric methods for post-processing, discuss the limitations of current approaches, and introduce COBASE, our new methodology.
+In Section 3, we review existing nonparametric and parametric methods for postprocessing, discuss the limitations of current approaches, and introduce COBASE, our new methodology.
 Section 4 tests COBASE in our case studies.
@@
 For this case study, we take operational forecasts of 2-meter temperature (T2m) and dew-point temperature (DPT) from the global ensemble prediction system run by the European Center for Medium-Range Weather Forecasting (ECMWF).
-We use the 51-members that are initialized each day at 12 UTC during the period 2021-2024, and take the forecasts that are valid at +24 hours.
+We use the 51-member ECMWF ensemble, consisting of one unperturbed control forecast and 50 perturbed ensemble members, that are initialized each day at 12 UTC during the period 2021-2024, and take the forecasts that are valid at +24 hours.
+The ensemble thus comprises two exchangeable groups: the statistically equivalent perturbed members and a distinct control member.
 There were several model upgrades during this time, including a change in horizontal resolution from  18km to  9km in 2023.
@@
 3 Methods
-This section outlines the two-step approaches currently considered state-of-the-art in multivariate statistical post-processing.
+This section outlines the two-step approaches currently considered state-of-the-art in multivariate statistical postprocessing.
 The key idea of these methods is to separate the correction of forecasts into two parts, one that addresses the marginal forecasts, and another that pertains the dependence structure.
-In practical terms, this involves initially applying univariate post-processing techniques to each margin separately, meaning that we address each weather variable, lead time, and station on an individual basis.
+In practical terms, this involves initially applying univariate postprocessing techniques to each margin separately, meaning that we address each weather variable, lead time, and station on an individual basis.
 Afterwards, we restore the missing dependencies by applying the theory of copulas.
@@
 Therefore, copulas can identify the dependence structure of a phenomenon without considering its marginal distributions.
-In the post-processing context, they can be used in the second step to reintroduce a dependence structure which gets lost in the univariate post-processing and sampling step.
+In the postprocessing context, they can be used in the second step to reintroduce a dependence structure which gets lost in the univariate postprocessing and sampling step.
 In the remainder of the section, we first introduce the univariate correction method used in our study, followed by the nonparametric and parametric copula-based approaches, and our newly proposed method COBASE.
 A summary of the specifications of the methods used with the corresponding labels is given in Table 2.
-3.1 Univariate Post-processing: Ensemble Model Output Statistics (EMOS)
+3.1 Univariate postprocessing: Ensemble Model Output Statistics (EMOS)
 One widely adopted method for univariate correction is Ensemble Model Output Statistics (EMOS) (Gneiting et al., 2005).
@@
 $\text{CRPS}(F_{ens},y)=\dfrac{1}{M}\sumop\displaylimits\limits_{m=1}^{M}|x_{m}-y|-\dfrac{1}{2M^{2}}\sumop\displaylimits\limits_{n=1}^{M}\sumop\displaylimits\limits_{m=1}^{M}|x_{n}-x_{m}|.$ (4)
-While EMOS provides effective univariate corrections, it does not capture dependencies across multiple variables or locations, motivating the use of multivariate post-processing methods.
-In the following section, we review non-parametric empirical copula-based postprocessing approaches.
-3.2 Nonparametric multivariate post-processing: Empirical-copula methods
-This section reviews the most common two-step nonparametric techniques for statistical post-processing.
+While EMOS provides effective univariate corrections, it does not capture dependencies across multiple variables or locations, motivating the use of multivariate postprocessing methods.
+In the following section, we review nonparametric empirical copula-based postprocessing approaches.
+3.2 Nonparametric multivariate postprocessing: Empirical-copula methods
+This section reviews the most common two-step nonparametric techniques for statistical postprocessing.
 All of these methods are based on empirical copulas, which offer an effective way to incorporate the dependence structure into weather forecasting problems without imposing any parametric assumptions (Schefzik, 2015).
@@
 $\bar{x}^{l,\tau}:=\dfrac{1}{M}\sumop\displaylimits_{m=1}^{M}x_{m}^{l,\tau},\quad s_{l,\tau}:=\sqrt{\dfrac{1}{M-1}\sumop\displaylimits_{m=1}^{M}\left(x_{m}^{l,\tau}-\bar{x}^{l,\tau}\right)^{2}}.$
-In addition to the construction of the reference dependence structure $DC_{ref}$, another point that can generally vary is the sampling technique used in S3 of Algorithm 1.
+In addition to the construction of the reference dependence structure $DC_{ref}$, another point that can generally vary is the sampling technique used in Step3 of Algorithm 1.
 In Schefzik et al.
@@
 Algorithm 1 Two-step rank-based postprocessing methods
-S1: Derive a reference dependence structure $DC_{ref}$;
-S2: Apply EMOS and obtain a corrected distribution $F_{l}$ for each margin $l=1,\ldots,d$;
+Step1: Derive a reference dependence structure $DC_{ref}$;
+Step2: Apply EMOS and obtain a corrected distribution $F_{l}$ for each margin $l=1,\ldots,d$;
 […]
-S4: Shuffle the corrected univariate samples according to the ranks of $DC_{ref}$.
+Step4: Shuffle the corrected univariate samples according to the ranks of $DC_{ref}$.
 In our case studies, we ensure a fair comparison with ECC by setting the size of the ensemble at $N=17$ for the ALADIN-LAEF groups and $N=51$ for the ECMWF data.
-For the Sim Schaake and Schaake Shuffle methods, we use all available observations and raw forecasts, including past and future dates (approximately four years), to ensure a consistent number of reference dates throughout the test set.
+For the Sim Schaake and Schaake Shuffle methods (in both datasets), we use all available observations and raw forecasts, including past and future dates after the validation date (approximately four years), to ensure a consistent number of reference dates throughout the test set.
+Although future observations are of course not available in an operational setting, we consider it reasonable to assume that four years of past observations and forecasts can be accessed, which is why we include future dates in this analysis.
 Specifically for the Schaake Shuffle, we select observations randomly from a 14-day window before and after the forecast date to preserve the seasonal climatology.
+We additionally examined the performance using larger window sizes and observed no notable differences in the outcomes.
 The main properties of the empirical copula-based approaches described here are summarized in Table 1. As emphasized in the table, the main limitations of these approaches are as follows.
   • (a)
-    The necessity for extensive historical data to more accurately illustrate the dependence influences both the Schaake Shuffle and Sim Schaake.
+    The direct reliance on past observations as dependence templates for reconstructing multivariate dependence, which affects both the Schaake Shuffle and Sim Schaake.
+This restricts the reconstructed dependence to patterns directly observed in the historical record and implicitly assumes that suitable historical analogues exist for the current forecast situation.
   • (b)
-    The requirement for extensive historical raw forecasts is crucial for the similarity criterion, impacting Sim Schaake.
+    The additional requirement of historical raw forecasts and a sufficiently large pool of past observations/forecasts to define a meaningful similarity criterion, which specifically impacts Sim Schaake.
   • (c)
-    The requirement that the sample size of the corrected forecasts ($N$) must equal the number of ensemble members ($M$), imposes a significant limitation on the applicability of ECC.
+    The requirement that the sample size of the corrected forecasts ($N$) must equal the number of ensemble members ($M$), which imposes a significant limitation on the applicability of ECC.
 Although these limitations might seem less relevant in an operational context, they may have a significant impact in practice.
-An example is Case Study 1 in Austria, where the post-processing of the operational system C-LAEF could not be performed due to the lack of historical raw forecasts and historical observations.
+An example is Case Study 1 in Austria, where the postprocessing of the operational system C-LAEF could not be performed due to the lack of historical raw forecasts and historical observations.
 Furthermore, for ALADIN-LAEF, the limited ensemble size of only $17$ members posed a substantial constraint on the application of the ECC method.
@@
 […]
-$\text{ES}=\dfrac{1}{M}\sumop\displaylimits\limits_{m=1}^{M}\|x_{m}-y\|-\dfrac{1}{2M^{2}}\sumop\displaylimits\limits_{\nu=1}^{M}\sumop\displaylimits\limits_{m=1}^{M}\|x_{\nu}-x_{m}\|$ (7)
+$\text{ES}=\dfrac{1}{M}\sumop\displaylimits\limits_{m=1}^{M}\|x_{m}-y\|-\dfrac{1}{2M^{2}}\sumop\displaylimits\limits_{\nu=1}^{M}\sumop\displaylimits\limits_{m=1}^{M}\|x_{\nu}-x_{m}\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}.}$ (7)
 In addition, the Variogram Score (VS) (Scheuerer and Hamill, 2015) explicitly addresses miscalibrations in the correlation structure of multivariate forecasts.
 In the ensemble setting, it is defined as
-$VS=\sumop\displaylimits_{i=1}^{d}\sumop\displaylimits_{j=1}^{d}{w}_{i,j}\left(\left|{y}_{i}-{y}_{j}\right|^{p}-\frac{1}{m}\sumop\displaylimits_{k=1}^{m}\left|x^{(k)}_{i}-x^{(k)}_{j}\right|^{p}\right)^{2}.$ (8)
+$VS=\sumop\displaylimits_{i=1}^{d}\sumop\displaylimits_{j=1}^{d}{w}_{i,j}\left(\left|{y}_{i}-{y}_{j}\right|^{p}-\frac{1}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}}\sumop\displaylimits_{k=1}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}}\left|x^{(k)}_{i}-x^{(k)}_{j}\right|^{p}\right)^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},}$ (8)
 where $w_{i,j}$ are weights and $p>0$ is the order of the score.
 Because the VS compares pairwise differences between variables in the observation and forecast, it is particularly sensitive to errors in the dependence structure.
-In our case study, we use both the Energy Score and the Variogram Score to assess and compare the performance of the different multivariate post-processing methods.
+In our case study, we use both the Energy Score and the Variogram Score to assess and compare the performance of the different multivariate postprocessing methods.
 These metrics provide complementary information, evaluating both the overall accuracy of the multivariate forecasts and their ability to represent dependence structures.
 We used the variant with no weights and $p=1$ for the Variogram Scores, but we obtained consistent results for other variants as well.
-3.3 Existing parametric multivariate post-processing
+Method Approach Dependence No direct use of past dependence templates No use of historical raw forecasts Flexible marginal sampling No restriction on the sample size
+Sim Schaake Shuffling Empirical copula (from past obs) × ×
+Schaake Shuffle Shuffling Empirical copula (from past obs) ×
+ECC Shuffling Empirical copula (from raw forecasts) ×
+Copula methods (parametric) Parametric fit (copula model) Parametric copula model (rolling window) ×
+COBASE Parametric fit (copula model) + shuffling Parametric copula model (rolling window)
+Table 1: Summary of the multivariate postprocessing methods analyzed in this work and their main characteristics.
+Check marks indicate that a method satisfies the corresponding property, while crosses indicate that it does not.
+3.3 Existing parametric multivariate postprocessing
 The reordering strategies discussed so far are, in fact, nonparametric.
@@
 Therefore, GCA can be considered as a genuine copula-based approach where $C$ is the Gaussian copula.
-The copula-based parametric correction steps for post-processing are summarized in Algorithm 2, which illustrates that any copula model can be used in place of the Gaussian copula.
-In our case studies, we estimate the copula parameters for standard parametric copula models using a 30-day rolling window, consistent with the EMOS procedure.
-We notice that there has been limited research on the application of alternative copulas in the post-processing context, with notable exceptions such as Jobst et al.
+The copula-based parametric correction steps for postprocessing are summarized in Algorithm 2, which illustrates that any copula model can be used in place of the Gaussian copula.
+In both case studies, copula parameters for the parametric copula models are estimated using a rolling 30-day training window.
+In initial experiments, longer training windows (e.g., 50 days) were also considered but did not lead to clear performance changes.
+Using longer windows increases the risk of mixing different seasonal statistics in the training data, while overly short windows may adversely affect the quality of dependence estimation.
+We therefore adopt a rolling 30-day window to ensure consistency with the EMOS-based univariate postprocessing, so that both univariate and multivariate corrections rely on the same pool of past observations.
+In our study, copula parameters are estimated on each rolling training window using pseudo-observations obtained from the marginal CDF values.
+For the Gaussian and Archimedean copulas considered here, the copula parameters are fitted using inversion of Kendall’s tau (Hofert et al., 2018), which provides a fast and robust moment-based estimator.
+Although parametric copula-based approaches also rely on past observations for parameter estimation, their dependence on historical data is less pronounced than for empirical shuffling methods such as the Schaake Shuffle or Sim Schaake, which directly use past observations as dependence templates.
+In parametric copula models, the estimated dependence structure can deviate from any individual historical realization, allowing for greater flexibility in representing multivariate dependence.
+We notice that there has been limited research on the application of alternative copulas in the postprocessing context, with notable exceptions such as Jobst et al.
 (2024) and Flos (2022).
@@
 Consequently, univariate skill often suffers even when multivariate dependencies are well captured.
-In the next section, we provide evidence that this issue is substantial and would affect the performance of any multivariate parametric post-processing method.
+In the next section, we provide evidence that this issue is substantial and would affect the performance of any multivariate parametric postprocessing method.
 Algorithm 2 Two-step copula-based postprocessing methods
-S1: Estimate the parameters $\alpha$ of a copula model $C_{\alpha}$;
-S2: Apply EMOS and obtain a corrected distribution $F_{l}$ for each margin $l=1,\ldots,d$;
+Step1: Estimate the parameters $\alpha$ of a copula model $C_{\alpha}$;
+Step2: Apply EMOS and obtain a corrected distribution $F_{l}$ for each margin $l=1,\ldots,d$;
 […]
-Figure 1: Boxplots of Diebold-Mariano test statistics for CRPS of T2m, comparing EMOS-Q against EMOS-R (baseline). Results are presented for the Austrian and Dutch stations used in this study.
+Figure 1: Boxplots of Diebold-Mariano test statistics Diebold and Mariano (1995) for CRPS of T2m, comparing EMOS-Q against EMOS-R (baseline).
+Results are presented for the Austrian and Dutch stations used in this study.
 Figure 2: Boxplots of Diebold-Mariano test statistics to compare ES and VS of Sim Schaake with uniform quantile sampling (SimSSh) versus Sim Schaake with randomly drawn margins (SimSSh-R, baseline).
 Results are presented for T2m for the three groups of Austrian stations and for T2m and DPT for the group of Dutch stations.
-3.3.1 Limitations of univariate random sampling in statistical post-processing
-As mentioned earlier, the reliance of some univariate and multivariate post-processing techniques on random sampling can lead to poorly calibrated margins.
+3.3.1 Limitations of univariate random sampling in statistical postprocessing
+As mentioned earlier, the reliance of some univariate and multivariate postprocessing techniques on random sampling can lead to poorly calibrated margins.
 To illustrate this, we compare the CRPS of T2m obtained for all stations in the study (Figure 1) using EMOS-R, i.e., random draws from the corrected Gaussian distribution, as the baseline.
 Figure 1 shows that drawing uniform quantiles instead (EMOS-Q) consistently outperforms EMOS-R across all stations.
-We further examine how univariate random sampling affects multivariate corrections by comparing the Energy Score and the Variogram Score for the Sim Schaake method with different marginal samples, i.e., EMOS-R versus EMOS-Q.
+We further examine how univariate random sampling may affect multivariate corrections.
+To illustrate this effect, we focus on the Sim Schaake method, which has been shown to perform best for the ALADIN-LAEF system (Perrone et al., 2020) and remains competitive for the ECMWF dataset considered here (Schefzik, 2016).
+We compare the Energy Score and the Variogram Score for Sim Schaake using different marginal sampling strategies, namely EMOS-R and EMOS-Q.
 Essentially, the two methods compared in Figure 2 share the same dependence structure and only differ in the marginal sampling technique.
@@
 This investigation highlights how random sampling of marginal distributions can negatively affect multivariate corrections.
-Since copula-based techniques fundamentally rely on random draws from parametric multivariate distributions, this limitation may systematically compromise their performance.
+Since copula-based techniques fundamentally rely on random draws from parametric multivariate distributions, this limitation may negatively affect their performance.
 To overcome this limitation, we develop a new method, COBASE, which is presented in the following section.
@@
 A random sample is then drawn from this copula to serve as the reference dependence structure $C_{ref}$ (Step S2).
-Next, standard univariate postprocessing, such as EMOS, is applied to each margin $l=1,\ldots,d$ to obtain corrected distributions $F_{l}$ (Step S3).
+Next, standard univariate postprocessing, such as EMOS, is applied to each margin $l=1,\ldots,d$ to obtain corrected distributions $F_{l}$ (Step3).
 Independent samples of size $N$ are drawn from these corrected margins using uniform quantiles (Step S4).
@@
 By doing so, COBASE preserves the dependence structure depicted by the parametric copula while leveraging the improved marginal calibration of advanced univariate sampling through, for example, uniform quantiles.
-As previously discussed, Table 1 provides an overview of the main multivariate post-processing methods and their key properties.
-In the table, COBASE stands out by combining the strengths of parametric and non-parametric two-step approaches: The parametric copula captures dependence, while uniform quantile sampling ensures well-calibrated margins.
+As previously discussed, Table 1 provides an overview of the main multivariate postprocessing methods and their key properties.
+In the table, COBASE stands out by combining the strengths of parametric and nonparametric two-step approaches: The parametric copula captures dependence, while uniform quantile sampling ensures well-calibrated margins.
 In the next section, we evaluate COBASE in two real-world case studies, comparing it with state-of-the-art methods to highlight its practical advantages.
@@
 Algorithm 3 Our proposed approach - COBASE
-S1: Estimate the parameters $\alpha$ of a copula model $C_{\alpha}$;
+Step1: Estimate the parameters $\alpha$ of a copula model $C_{\alpha}$;
 […]
-S3: Apply EMOS and obtain a corrected distribution $F_{l}$ for each margin $l=1,\ldots,d$;
+Step3: Apply EMOS and obtain a corrected distribution $F_{l}$ for each margin $l=1,\ldots,d$;
 […]
-S5: Shuffle the corrected univariate samples according to the ranks of $C_{ref}$.
-Method Approach Dependence Free from long historical data Free from historical raw forecasts Flexible marginal sampling No restriction on the sample Size
-Sim Schaake Shuffling Empirical copula (from past obs) × ×
-Schaake shuffle Shuffling Empirical copula (from past obs) ×
-ECC Shuffling Empirical copula (from raw forecasts) ×
-Copula methods (parametric) Parametric fit (copula model) Parametric copula model (rolling window) ×
-COBASE Parametric fit (copula model) + shuffling Parametric copula model (rolling window)
-Table 1: Summary of the multivariate post-processing methods analyzed in this work, alongside their main characteristics.
+Step5: Shuffle the corrected univariate samples according to the ranks of $C_{ref}$.
 Figure 3: Boxplots of Diebold-Mariano test statistics for the case studies.
@@
 4 Results
-We evaluate the proposed COBASE method in comparison with state-of-the-art post-processing approaches on the two real-world case studies described in Section 2.
+We evaluate the proposed COBASE method in comparison with state-of-the-art postprocessing approaches on the two real-world case studies described in Section 2.
 Although the focus of this work is not on univariate correction, CRPS results for 2-meter temperature across all stations are included in Table 3 for completeness.
-These results show that EMOS-Q, which uses uniform quantile sampling, consistently outperforms methods based on random sampling, clearly indicating that quantile-based approaches are generally preferable.
+These results show that EMOS-Q, which uses uniform quantile sampling, outperforms methods based on random sampling on average, indicating that quantile-based approaches are generally preferable in this setting.
+The reported differences are based on mean CRPS values computed over the full evaluation period.
+Although averaging may mask the influence of individual validation days, the overall results consistently favor quantile-based sampling.
+Small differences in mean CRPS among methods relying on random sampling of the marginal distributions may partly reflect the variability induced by the random draws, which is a known limitation of EMOS-R and traditional copula-based approaches.
+This source of variability is avoided by the deterministic sampling method used in EMOS-Q.
 The main focus of this section is on multivariate forecast performance, assessed using the Energy Score (ES) and the Variogram Score (VS).
@@
 5 Discussion and conclusion
-This paper introduces COBASE, a new multivariate post-processing method that combines the structure of parametric copula models with the flexibility of non-parametric methods.
-By decoupling the sampling of margins from the copula-based dependence structure, COBASE addresses a key issue in the copula-based approaches for post-processing, namely, the degradation in forecast quality caused by random sampling from marginal distributions.
+This paper introduces COBASE, a new multivariate postprocessing method that combines the structure of parametric copula models with the flexibility of nonparametric methods.
+By decoupling the sampling of margins from the copula-based dependence structure, COBASE addresses a key issue in the copula-based approaches for postprocessing, namely, the degradation in forecast quality caused by random sampling from marginal distributions.
 A key strength of COBASE is its flexibility.
@@
 Our results from two real-world case studies in Austria and the Netherlands show that COBASE consistently improves upon standard copula-based methods, including GCA, and outperforms Schaake Shuffle.
-Moreover, COBASE achieves performance comparable to widely used non-parametric techniques such as Sim Schaake and ECC, while avoiding their structural limitations and extensive data requirements.
+Moreover, COBASE achieves performance comparable to widely used nonparametric techniques such as Sim Schaake and ECC, while avoiding their structural limitations and extensive data requirements.
 A closer look at the scores in Table 4 reveals that COBASE-GCA performs similarly to Sim Schaake, even in the station groups where the latter shows a slight advantage.
@@
 This shorter training horizon makes COBASE substantially more flexible and easier to apply in operational contexts.
-For example, COBASE would be applicable to C-LAEF, while Sim Schaake would not be suitable in that case due to the lack of past raw forecasts.
+For example, COBASE would be applicable to C-LAEF, whereas Sim Schaake would not be suitable in that case due to the lack of past raw forecasts.
+While Lakatos et al.
+(2023) report that increasing the number of historical pool dates for the Schaake Shuffle does not necessarily lead to improved forecast performance, the Schaake Shuffle still relies on historical dependence templates and exhibits weaker average performance in our experiments.
+In contrast, COBASE achieves improved skill on average while requiring only a short rolling training window, making it a more practical choice in operational settings.
 Additionally, COBASE shows robust performance even in relatively high-dimensional settings, such as the ECMWF ENS group with 12 variables, and no dimension-specific tuning.
+Another important aspect is that COBASE, as any parametric copula-based method, can be applied to both exchangeable and nonexchangeable ensemble systems, whereas this may not necessarily hold for ECC and Sim Schaake, depending on the underlying assumptions and, in the latter case, on the chosen similarity criterion.
+While ECC formally relies on ensemble exchangeability, previous work in Perrone et al.
+(2020) has shown that nonexchangeability does not necessarily prevent its application, although it requires careful diagnostic analysis to identify potential anomalous dependence structures in the raw forecasts.
+Such considerations depend on the specific characteristics of the NWP system and are therefore model-dependent.
+For this reason, we include ECC in this study for comparison purposes.
 The added value of COBASE becomes particularly evident in comparison to that of its corresponding standard copula-based variant.
 In this case, the performance gap is much more pronounced, confirming that the observed improvements stem from the structure of the COBASE method itself rather than from the specific copula family employed.
-This supports the main goal of the study: Not to identify the single best-performing copula, but to demonstrate that, across different configurations, COBASE consistently enhances the effectiveness of copula-based post-processing.
+This supports the main goal of the study: Not to identify the single best-performing copula, but to demonstrate that, across different configurations, COBASE consistently enhances the effectiveness of copula-based postprocessing.
 In this sense, COBASE represents a more efficient, flexible, and broadly applicable alternative to conventional parametric multivariate approaches.
@@
 However, implementing dynamically varying copulas would also increase computational demands, so developing efficient strategies for adaptive copula selection or updating will be essential for practical applications.
-Finally, while COBASE avoids the limitations of random marginal sampling, it still relies on random draws from the fitted copula to represent the multivariate structure.
+Then, while COBASE avoids the limitations of random marginal sampling, it still relies on random draws from the fitted copula to represent the multivariate structure.
 Future research could explore alternative multivariate sampling strategies to reduce sampling variability and better capture the characteristics of the dependence structure, especially in the tails of the multivariate distributions.
 These avenues offer promising directions for further strengthening COBASE’s performance across a wider range of postprocessing tasks.
-To conclude, taken together, our findings demonstrate that COBASE already provides a simple, flexible, and effective framework for multivariate ensemble post-processing.
-By combining the strengths of parametric and non-parametric approaches, it improves forecast skill without requiring long historical records or raw ensemble data.
-The COBASE methodology is also particularly well suited for research on post-processing of multivariate dependence, as it allows a clear separation between the evaluation of the dependence structure and the marginals, thereby facilitating a better identification of potential improvements in the dependence component itself.
+Finally, an important direction for future work is the development of controlled simulation studies to further assess the performance of COBASE relative to benchmark methods, and to systematically investigate the role of ensemble size and ensemble exchangeability.
+To conclude, taken together, our findings demonstrate that COBASE already provides a simple, flexible, and effective framework for multivariate ensemble postprocessing.
+By combining the strengths of parametric and nonparametric approaches, it improves forecast skill without requiring long historical records or raw ensemble data.
+The COBASE methodology is also particularly well suited for research on postprocessing of multivariate dependence, as it allows a clear separation between the evaluation of the dependence structure and the marginals, thereby facilitating a better identification of potential improvements in the dependence component itself.
 These properties make COBASE particularly well-suited for operational settings with limited data or for newly developed forecasting systems, such as the C-LAEF system in Austria.
 Building on this foundation, future work will focus on adjusting the method to higher-dimensional settings, improving copula selection strategies, and enhancing the multivariate sampling step.
-With these developments, COBASE has the potential to become a robust and scalable solution for a wide range of ensemble post-processing applications.
+With these developments, COBASE has the potential to become a robust and scalable solution for a wide range of ensemble postprocessing applications.
+Acknowledgment
+We thank the two anonymous reviewers and the associate editor for their thoughtful comments, which helped improve an earlier version of the manuscript.
+Author Contribution Indication
+Maurits Flos: Conceptualization, Data curation, Formal analysis, Investigation, Software, Validation
+Bastien François: Conceptualization, Data curation, Investigation, Validation, Visualization, Writing – original draft, Writing – review & editing
+Irene Schicker: Data curation, Writing – original draft, Writing – review & editing
+Kirien Whan: Conceptualization, Data curation, Investigation, Software, Validation, Visualization, Writing – original draft, Writing – review & editing
+Elisa Perrone: Conceptualization, Formal analysis, Investigation, Methodology, Software, Supervision, Validation, Visualization, Writing – original draft, Writing – review & editing
 Data availability statement
-The datasets analyzed during this study are subject to restrictions and are therefore not publicly available.
+The data that support the findings of this study are available from the corresponding author upon reasonable request.
 To facilitate reproducibility, all analysis scripts and mock data that closely replicate the original datasets, allowing reproduction of the figures and tables, are available at https://github.com/elisaperrone/COBASE_github.
+Conflict of interest statement
+The authors have no conflicts of interest to declare.
 Appendix A Abbreviations
@@
 COBASE - Gumbel Copula + EMOS-Q COBASE-Gumbel
-Table 2: Overview of the statistical post-processing methods evaluated in the study, grouped by approach type, with their corresponding labels.
+Table 2: Overview of the statistical postprocessing methods evaluated in the study, grouped by approach type, with their corresponding labels.
 Appendix B Scores